# On Discrete Morse-Bott Theory

### v1

@@ Abstract.
This paper shows that discrete Morse–Bott theory can be developed in an intuitive way, which is achieved by improving the definition of the discrete Morse–Bott function originally introduced by S. Yaptieu. In fact, we demonstrate that various natural properties hold for discrete Morse–Bott functions and, in particular, establish the discrete Morse–Bott inequality, which can be regarded as an extension of both the discrete Morse inequalities and the continuous Morse–Bott inequalities.
@@ 1. Introduction
Morse theory has been developed as a framework to describe the topological structure of manifolds through the critical points of smooth functions (cf. [milnor1963morse]). It establishes a correspondence between the properties of critical points of Morse functions and the topology of manifolds. R. Bott [bott1954MBT] generalized Morse theory to Morse–Bott theory by allowing isolated critical points to be replaced by critical submanifolds. Morse–Bott theory is useful for studying functions with symmetries or conserved quantities and has applications in symplectic geometry (cf. [austin1995morse, banyaga2004lectures]).
R. Forman [forman1998DMT] introduced discrete Morse theory, which is a combinatorial version of Morse theory defined on CW complexes. Thus, it is suitable for computing homology and simplifying cell complexes, making it useful in various fields such as topological data analysis [harker2014discrete, mischaikow2013morse].
S. Yaptieu [yaptieu2017discrete] proposed a generalized version of discrete Morse theory, called a discrete Morse–Bott function, where isolated critical points are replaced by more general subsets. This generalization reflects the structure of Morse–Bott theory in a discrete setting and allows the study of discrete functions with symmetry. However, the definition in the preprint slightly differs from intuitive expectations, which implies the existence of some gaps (e.g. Remark 2.1(ii), the proofs of Lemmas 2.3, 2.4, 2.8). Therefore, we improve the definition to fill these gaps and make it more intuitive by extending the concept of critical points in discrete Morse theory to allow critical sets. In fact, we demonstrate a discrete Morse–Bott inequality, which naturally generalizes the discrete Morse inequalities known from Morse theory and can be derived from the continuous Morse–Bott theory. The inequality gives an upper bound for the Betti numbers of the underlying complex based on the structure of the critical sets. Furthermore, it opens the way for the analysis of data with group actions or symmetries.
The present paper consists of four sections. In the next section, we review the fundamental concepts and properties of discrete Morse theory. In § 3, we introduce discrete Morse-Bott function, which is a generalization of discrete Morse function (Theorem 3.2). In the final section, we observe that various properties that hold for discrete Morse functions also hold for discrete Morse–Bott functions, and establish the discrete Morse–Bott inequality (Theorem 4.12), which can be regarded as an extension of the discrete Morse inequalities (Proposition 4.13).
@@ 2.1. Fundamental concepts
Let $X,Y$ be topological spaces and $f\colon A\to X$ a continuous map on a subspace $A\subseteq Y$. Then we define
@@
where the equivalence relation $\sim$ is the smallest one such that for every $a\in A$, we have $a\sim f(a)$. The map $f$ is called an attaching map.
@@ A CW complex $X$ is constructed inductively as follows:
By definition, note that we have $X=\coprod_{\lambda\in\Lambda}e_{\lambda}^{n}$.
For simplicity, we identify a CW complex $\mathbb{K}$ with the family of all its cells. With a slight abuse of notation, we also use $\mathbb{K}$ to denote this family. From now on, let $\mathbb{K}$ be a CW complex. We have the following concepts and notations.
@@
A $k$-dimenonal face $\sigma$ of $\tau$ is regular (or $\sigma\overset{\mathrm{reg}}{{<}}\tau$) if the following conditions hold:
@@
(2) $\overline{\varphi^{-1}_{\tau}(\sigma)}$ is a closed $k$-ball.
@@
A cell $\sigma$ is a irregular face of $\tau$ (or $\sigma\overset{\mathrm{irr}}{{<}}\tau$) if $\sigma{<}\tau$ but $\sigma\overset{\mathrm{reg}}{\not<}\tau$.
@@
For a function on a CW complex, we call a cell nonritical when the value is not less than that on an adjacent higher-dimensional cell, as follows.
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a function defined on a CW complex $\mathbb{K}$. A cell $\sigma$ is a noncritical face of $\tau$ (or $\sigma\overset{\mathrm{nc}}{\prec}\tau$) if $\sigma\prec\tau\,$ and $f(\sigma)\geq f(\tau)$.
@@
Note that the concept of “noncritical” is also used in the definition of a discrete Morse–Bott function below.
We define the following “interval”.
@@ 2.2. Fundamental properties
Assume that $\alpha\overset{\mathrm{reg}}{<}\tau$. If $\dim\tau-\dim\alpha\geq 2$, then Lemma 2.2 implies that there is a cell $\beta\in\mathbb{K}$ such that $\alpha<\beta<\tau$, which contradicts the uniqueness of $\alpha$. Thus $\dim\tau-\dim\alpha=1$. However, Lemma 2.3 implies the existence of another cells between $\nu$ and $\tau$, which contradicts the uniqueness of $\alpha$. ∎
@@
In the previous lemma, the statement $\nu\overset{\mathrm{irr}}{\prec}\alpha$ need not hold. In fact, there is a CW complex satisfies $X=\nu\cup\widetilde{\nu}\cup\alpha\cup\tau$, $\nu\overset{\mathrm{reg}}{\prec}\alpha$ as in Figure 1.
@@ 2.3. Discrete Morse function
The function $f$ is discrete Morse if the following conditions holds for any $\sigma^{(k)}\in\mathbb{K}$:
(M1) For any cell $\tau$ with $\sigma\overset{\mathrm{irr}}{\prec}\tau$, we have $f(\sigma)<f(\tau)$.
@@
(M3) For any cell $\nu$ with $\nu\overset{\mathrm{irr}}{\prec}\sigma$, we have $f(\nu)<f(\sigma)$.
@@
From the definition (M1) of discrete Morse function, it follows that, for any discrete Morse function $f\colon\mathbb{K}\to\mathbb{R}$ and any cells $\sigma,\tau\in\mathbb{K}$, the condition $\sigma\overset{\mathrm{nc}}{\prec}\tau$ implies $\sigma\overset{\mathrm{reg}}{\prec}\tau$.
Lemma 2.5 (Lemma 6.11.[knudson2015morse]).
@@ 3.1. Definition of discrete Morse-Bott function
A collection $C$ for $f$ is a maximal family of cells satisfying the following conditions:
@@
Denote by $\mathcal{C}_{f}$ the set of collections of $f$. Then $\mathbb{K}=\bigsqcup\mathcal{C}_{f}$. For any cell $\sigma\in\mathbb{K}$, denote by $\mathcal{C}_{f}(\sigma)$ the collection containing $\sigma$.
Notice that the condition (2) in the previous definition is used only in Remark 3, Theorem 3.2, and Proposition 4.13.
For any collection $C$ for $f$ and any $k$-dimensional cell $\sigma^{(k)}\in C$, set $U^{C}(\sigma)$ and $D^{C}(\sigma)$ as follows:
$\begin{split}U^{C}(\sigma):=\#\{\tau^{(k+1)}\notin C\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}=&\#\{\tau^{(k+1)}\notin C\mid\sigma\prec\tau,f(\sigma)\geq f(\tau)\}\end{split}$
$D^{C}(\sigma):=\#\{\nu^{(k-1)}\notin C\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}=\#\{\nu^{(k-1)}\notin C\mid\nu\prec\sigma,f(\nu)\geq f(\sigma)\}$
Definition 20 (Discrete Morse-Bott function).
Let $\mathcal{C}:=\{C^{1},\ldots,C^{l}\}$ be the family of collections for a function $f\colon\mathbb{K}\to\mathbb{R}$ on a CW complex $\mathbb{K}$. The function $f$ is a discrete Morse-Bott function (with respect to $\mathcal{C}$) if the following conditions holds for any $i\in\{1,\ldots,l\}$ and for all $\sigma^{(k)}\in C^{i}$:
(MB1) For any cell $\tau$ with $\sigma\overset{\mathrm{irr}}{<}\tau$, we have $f(\sigma)<f(\tau)$ (and so $\tau\notin C^{i}$).
(MB2) $U^{C^{i}}(\sigma)\leq 1$.
(MB3) For any cell $\nu$ with $\nu\overset{\mathrm{irr}}{<}\sigma$, we have $f(\nu)<f(\sigma)$ (and so $\nu\notin C^{i}$).
(MB4) $D^{C^{i}}(\sigma)\leq 1$.
Example 2.
The function on the left in the figure is a discrete Morse–Bott function. On the other hand, the function on the right does not, since it fails to satisfy (MB4) at a one-dimensional cell $\sigma$. Indeed, we obtain $U^{\mathcal{C}_{f}(\sigma)}(\sigma)=2$, which contradicts (M4).
[…]
We have the following observation.
Lemma 3.1.
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse function. The function $f$ is a discrete Morse-Bott function such that $C\leq 2$ for any collection $C\in\mathcal{C}_{f}$.
Proof.
Fix any $\sigma\in\mathbb{K}$. The condition (MB1) (resp. (MB3)) follows from the condition (B1) (resp. (B3)). The following conditions
$U^{\mathcal{C}_{f}(\sigma)}(\sigma)\leq U(\sigma)\leq 1$
$D^{\mathcal{C}_{f}(\sigma)}(\sigma)\leq D(\sigma)\leq 1$
imply the conditions (MB2) and (MB4). This means that $f$ is a discrete Morse-Bott function.
Assume that $C\geq 3$. Then there is a $c$-path $\sigma_{0},\sigma_{1},\sigma_{2}$. Therefore, we obtain that $U(\sigma_{1})+D(\sigma_{1})\geq 2$, which contradicts $U(\sigma_{1})+D(\sigma_{1})=1$ because of Lemma 2.5. Thus $C\leq 2$. ∎
3.2. Reduced collection
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function. Recall that a cell $\sigma\in\mathbb{K}$ is critical if and only if $U(\sigma)=D(\sigma)=0$. Fix any collection $C\in\mathcal{C}_{f}$. We introduce the following concepts.
Definition 21.
A cell $\sigma\in C$ is upward noncritical if $U^{C}(\sigma)=1$.
Definition 22.
A cell $\sigma\in C$ is downward noncritical if $D^{C}(\sigma)=1$.
Definition 23 (Reduced collection).
The reduced collection $C^{\mathrm{red}}$ of $C$ is defined as the set obtained from $C$ by deleting all upward and downward noncritical cells in $C$.
@@ 3.2. Reduced collection
(2) $C=C^{\mathrm{red}}=\{\sigma\}$, where $C\in\mathcal{C}_{f}$ is the collection containing $\sigma$.
@@ 3.3. Generalization of a discrete Morse function
We have the following equivalence, which demonstrates that the discrete Morse–Bott function is a generalization of the discrete Morse function.
@@
(2) The function $f$ is discrete Morse-Bott such that, for any collection $C\in\mathcal{C}_{f}$, either #C=1\#C=1 or both $C=C^{\mathrm{red}}$ and #C=2\#C=2.
@@
Suppose that $f$ is discrete Morse. By Proposition 3.1, we have that $C\leq 2$ for any collection $C\in\mathcal{C}_{f}$. If #C=2\#C=2, then Lemma 2.5 implies that $0=U^{C}(\sigma)+D^{C}(\sigma)$ and so that $C=C^{\mathrm{red}}$, which implies assertion (2).
Conversely, suppose that assertion (2) holds. Fix any collection $C\in\mathcal{C}_{f}$ and any cell $\sigma\in C$. If $C$ is a singleton (i.e. $C=\{\sigma\}$), then $U(\sigma)=U^{C}(\sigma)$ and $D(\sigma)=D^{C}(\sigma)$ and so (MB2) (resp. (MB4)) for $C$ implies (B2) (resp. (B4)) for $C$. If #C=2\#C=2, then
$U(\sigma)=\#\{\tau\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}$ $=\#\{\tau\notin C\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}+\#\{\tau\in C\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}$
@@
$D(\sigma)=\#\{\tau\mid\tau\overset{\mathrm{nc}}{\prec}\sigma\}$ $=\#\{\tau\notin C\mid\tau\overset{\mathrm{nc}}{\prec}\sigma\}+\#\{\tau\in C\mid\tau\overset{\mathrm{nc}}{\prec}\sigma\}$
@@
and so (B2) and (B4) hold. This means that $f$ is discrete Morse. ∎

### v2

@@ Abstract.
This paper shows that discrete Morse-Bott theory can be developed as a natural extension of R. Forman’s discrete Morse theory by improving the definition of the discrete Morse-Bott function introduced by S. Yaptieu. To this end, we demonstrate that the combinatorial structure of critical cells can be extended to critical sets intuitively. Furthermore, we establish the discrete Morse-Bott inequalities, providing a unified view that extends both the discrete Morse inequalities and the continuous Morse-Bott inequalities.
@@ 1. Introduction
Morse theory has been developed as a framework for describing the topological structure of manifolds via the critical points of smooth functions (cf. [10]). It establishes a correspondence between the properties of critical points of Morse functions and the topology of manifolds. R. Bott [5] generalized Morse theory to Morse-Bott theory by allowing isolated critical points to be replaced by critical submanifolds. Morse-Bott theory is useful for studying functions with symmetries or conserved quantities and has applications in symplectic geometry (cf. [1, 2]).
R. Forman [7] introduced discrete Morse theory, which is a combinatorial version of Morse theory defined on CW complexes. Thus, it is suitable for computing homology and simplifying cell complexes, making it useful in various fields such as topological data analysis [8, 11].
S. Yaptieu [12] proposed a generalized version of discrete Morse theory, called a discrete Morse-Bott function, where isolated critical points are replaced by more general subsets. This generalization reflects the structure of Morse-Bott theory in a discrete setting and allows the study of discrete functions with symmetry. However, although this work remains a preprint from 2017, the definition slightly differs from intuitive expectations, which implies the existence of some gaps (e.g. Remark 2.1(ii), the proofs of Lemmas 2.3, 2.4, 2.8). Therefore, we improve the definition to fill these gaps and make it more intuitive by extending the concept of critical points in discrete Morse theory to allow critical sets. In fact, we demonstrate a discrete Morse-Bott inequality, which naturally generalizes the discrete Morse inequalities known from the discrete Morse theory [6, 7] and can be derived from the continuous Morse-Bott theory [4]. The inequality gives an upper bound for the Betti numbers of the underlying complex based on the structure of the critical sets. Furthermore, it enables analysis of data using group actions or symmetries.
The present paper consists of four sections. In the next section, we review the fundamental concepts and properties of discrete Morse theory. In § 3, we introduce a discrete Morse-Bott function, which is a generalization of a discrete Morse function (Theorem 3.2). In the final section, we observe that various properties that hold for discrete Morse functions also hold for discrete Morse-Bott functions, and establish the discrete Morse-Bott inequality (Theorem 4.12), which can be regarded as an extension of the discrete Morse inequalities (Proposition 4.13).
@@ 2.1. Fundamental concepts
Let $X,Y$ be topological spaces and $f\colon A\to X$ a continuous map on a subspace $A\subseteq Y$. Then we define a subset
@@
where $\sqcup$ denotes a disjoint union and the equivalence relation $\sim$ is the smallest one such that for every $a\in A$, we have $a\sim f(a)$. The map $f$ is called an attaching map.
@@ A CW complex $X$ is constructed inductively as follows:
By definition, note that we have $X=\coprod_{\lambda\in\Lambda}e_{\lambda}^{n}$. For simplicity, we identify a CW complex $\mathbb{K}$ with the family of all its cells. With a slight abuse of notation, we also use $\mathbb{K}$ to denote this family. From now on, let $\mathbb{K}$ be a CW complex. We have the following concepts and notations.
@@
A $k$-dimensional face $\sigma$ of $\tau$ is regular (or $\sigma\overset{\mathrm{reg}}{{<}}\tau$) if the following conditions hold:
@@
(2) $\overline{\varphi^{-1}_{\tau}(\sigma)}$ is a $k$-dimensional closed disk.
@@
A cell $\sigma$ is an irregular face of $\tau$ (or $\sigma\overset{\mathrm{irr}}{{<}}\tau$) if $\sigma{<}\tau$ but $\sigma\overset{\mathrm{reg}}{\not<}\tau$.
@@
For a function on a CW complex, we call a cell noncritical when the value is not less than that on an adjacent higher-dimensional cell, as follows.
@@
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a function defined on a CW complex $\mathbb{K}$. A cell $\sigma$ is a noncritical facet of $\tau$ (or $\sigma\overset{\mathrm{nc}}{\prec}\tau$) if $\sigma\prec\tau\,$ and $f(\sigma)\geq f(\tau)$.
@@
Note that the concept of “noncritical” is also used in the definition of a discrete Morse-Bott function below. We define the following “interval”.
@@ 2.2. Fundamental properties
Assume that $\alpha\overset{\mathrm{reg}}{<}\tau$. If $\dim\tau-\dim\alpha\geq 2$, then Lemma 2.2 implies that there is a cell $\beta\in\mathbb{K}$ such that $\alpha<\beta<\tau$, which contradicts the uniqueness of $\alpha$. Thus $\dim\tau-\dim\alpha=1$. However, Lemma 2.3 implies the existence of another cell between $\nu$ and $\tau$, which contradicts the uniqueness of $\alpha$. ∎
@@
In the previous lemma, the statement $\nu\overset{\mathrm{irr}}{\prec}\alpha$ need not hold. In fact, there is a CW complex that satisfies $X=\nu\cup\widetilde{\nu}\cup\alpha\cup\tau$, $\nu\overset{\mathrm{reg}}{\prec}\alpha$ as in Figure 1.
@@ 2.3. Discrete Morse function
The function $f$ is discrete Morse [7] if the following conditions hold for any $\sigma\in\mathbb{K}$:
(M1) For any cell $\tau$ with $\sigma\overset{\mathrm{irr}}{<}\tau$, we have $f(\sigma)<f(\tau)$.
@@
(M3) For any cell $\nu$ with $\nu\overset{\mathrm{irr}}{<}\sigma$, we have $f(\nu)<f(\sigma)$.
@@
From definition (M1) of discrete Morse function, it follows that, for any discrete Morse function $f\colon\mathbb{K}\to\mathbb{R}$ and any cells $\sigma,\tau\in\mathbb{K}$, the condition $\sigma\overset{\mathrm{nc}}{\prec}\tau$ implies the condition $\sigma\overset{\mathrm{reg}}{\prec}\tau$.
Lemma 2.5 (Lemma 6.11.[9]).
@@ 2.4. Fundamental concepts
We recall the combinatorial vector field introduced by Forman [6] as follows.
@@ 3.1. Definition of discrete Morse-Bott function
For any $k$-dimensional cell $\sigma^{(k)}$, set $U^{\mathrm{snc}}(\sigma)$ and $D^{\mathrm{snc}}(\sigma)$ as follows:
$\begin{split}U^{\mathrm{snc}}(\sigma):=\#\{\tau^{(k+1)}\mid\sigma\overset{\mathrm{snc}}{\prec}\tau\}=\#\{\tau^{(k+1)}\mid\sigma\prec\tau,f(\sigma)>f(\tau)\}\end{split}$
$\begin{split}D^{\mathrm{snc}}(\sigma):=\#\{\nu^{(k-1)}\mid\nu\overset{\mathrm{snc}}{\prec}\sigma\}=\#\{\nu^{(k-1)}\mid\nu\prec\sigma,f(\nu)>f(\sigma)\}\end{split}$
Definition 18 (Discrete Morse-Bott function).
A function $f\colon\mathbb{K}\to\mathbb{R}$ on a CW complex $\mathbb{K}$ is a discrete Morse-Bott function if the following conditions hold for any $\sigma^{(k)}\in\mathbb{K}$:
(M1) For any cell $\tau$ with $\sigma\overset{\mathrm{irr}}{<}\tau$, we have $f(\sigma)<f(\tau)$.
(MB2) $U^{\mathrm{snc}}(\sigma)\leq 1$.
(M3) For any cell $\nu$ with $\nu\overset{\mathrm{irr}}{<}\sigma$, we have $f(\nu)<f(\sigma)$.
(MB4) $D^{\mathrm{snc}}(\sigma)\leq 1$.
3.1.1. Collections and paths
@@
A collection $L$ for $f$ is a maximal family of cells satisfying the following conditions:
(1) There is a number $c\in\mathbb{R}$ such that $L\subset f^{-1}(c)$.
(2) For any $\sigma,\tau\in L$, there is a $c$-path from $\sigma$ to $\tau$.
Notice that any collection for $f$ coincides with an equivalence class for the equivalence relation given by the concept of $r$-path for some $r\in\mathbb{R}$. Denote by $\mathcal{L}_{f}$ the set of collections of $f$. Then $\mathbb{K}=\bigsqcup\mathcal{L}_{f}$. For any cell $\sigma\in\mathbb{K}$, denote by $\mathcal{L}_{f}(\sigma)$ the collection containing $\sigma$.
Notice that condition (2) in the previous definition is used only in Remark 4, Theorem 3.2, and Proposition 4.13.
For any collection $L$ for $f$ and any $k$-dimensional cell $\sigma^{(k)}\in L$, the following relations hold :
$\begin{split}U^{\mathrm{snc}}(\sigma)=\#\{\tau^{(k+1)}\notin L\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}&=\#\{\tau^{(k+1)}\notin L\mid\sigma\prec\tau,f(\sigma)\geq f(\tau)\}\end{split}$
$\begin{split}D^{\mathrm{snc}}(\sigma)=\#\{\nu^{(k-1)}\notin L\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}&=\#\{\nu^{(k-1)}\notin L\mid\nu\prec\sigma,f(\nu)\geq f(\sigma)\}\end{split}$
Example 2.
The function on the left in Figure 2 is a discrete Morse-Bott function. On the other hand, the function on the right in Figure 2 is not, since it fails to satisfy (MB4) at a one-dimensional cell $\sigma$. Indeed, we obtain $U^{\mathrm{snc}}(\sigma)=2$, which contradicts (M4).
[…]
Figure 2. Left: discrete Morse-Bott function; right: non-discrete Morse-Bott function.
We have the following observation.
Lemma 3.1.
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse function. The function $f$ is a discrete Morse-Bott function such that $L\leq 2$ for any collection $L\in\mathcal{L}_{f}$.
Proof.
Fix any $\sigma\in\mathbb{K}$. conditions (M1) and (M3) follow from definition of a discrete Morse function. The following conditions
$U^{\mathrm{snc}}(\sigma)\leq U(\sigma)\leq 1$
$D^{\mathrm{snc}}(\sigma)\leq D(\sigma)\leq 1$
imply conditions (MB2) and (MB4). This means that $f$ is a discrete Morse-Bott function.
Assume that $L\geq 3$. Then there is a $c$-path $\sigma_{0},\sigma_{1},\sigma_{2}$. Therefore, we obtain that $U(\sigma_{1})+D(\sigma_{1})\geq 2$, which contradicts $U(\sigma_{1})+D(\sigma_{1})=1$ because of Lemma 2.5. Thus $L\leq 2$. ∎
3.2. Reduced collection
Let $f\colon\mathbb{K}\to\mathbb{R}$ be a discrete Morse-Bott function. Recall that a cell $\sigma\in\mathbb{K}$ is critical if and only if $U(\sigma)=D(\sigma)=0$. We introduce the following concepts.
Definition 21.
A cell $\sigma\in\mathbb{K}$ is upward noncritical if $U^{\mathrm{snc}}(\sigma)=1$.
Definition 22.
A cell $\sigma\in\mathbb{K}$ is downward noncritical if $D^{\mathrm{snc}}(\sigma)=1$.
Definition 23.
A cell $\sigma\in\mathbb{K}$ is weakly critical if $D^{\mathrm{snc}}(\sigma)=U^{\mathrm{snc}}(\sigma)=0$.
Definition 24 (Reduced collection).
A reduced collection $C$ for $f$ is a maximal family of cells satisfying the following conditions:
@@ 3.2. Reduced collection
(3) For any $\sigma\in C$, $\sigma$ is weakly critical.
Denote by $\mathcal{C}_{f}$ the set of reduced collections of $f$. For any cell $\sigma\in\mathbb{K}$, denote by $\mathcal{C}_{f}(\sigma)$ the reduced collection containing $\sigma$.
Remark 3.
For any reduced collection $C\in\mathcal{C}_{f}$, there exists a unique collection $L$ such that $C\subset L$.
Denote by $L(C)$ the collection containing $C$.
@@
(2) $L(\mathcal{C}_{f}(\sigma))=\mathcal{C}_{f}(\sigma)$.
@@ 3.3. Generalization of a discrete Morse function
We have the following equivalence, which demonstrates that the discrete Morse-Bott function is a generalization of the discrete Morse function.
@@
(2) The function $f$ is discrete Morse-Bott such that, for any collection $L\in\mathcal{L}_{f}$, one of the following holds:
(i) #L=1\#L=1.
(ii) #L=2\#L=2 and $L$ is a reduced collection.
@@
Suppose that $f$ is discrete Morse. By Proposition 3.1, we have that $L\leq 2$ for any collection $L\in\mathcal{L}_{f}$. If #L=2\#L=2, then Lemma 2.5 implies that $0=U^{\mathrm{snc}}(\sigma)+D^{\mathrm{snc}}(\sigma)$ and so that $L$ is a reduced collection, which implies assertion (2).
Conversely, suppose that assertion (2) holds. Fix any collection $L\in\mathcal{L}_{f}$ and any cell $\sigma\in L$. If $L$ is a singleton (i.e. $L=\{\sigma\}$), then $U(\sigma)=U^{\mathrm{snc}}(\sigma)$ and $D(\sigma)=D^{\mathrm{snc}}(\sigma)$ and so (MB2) (resp. (MB4)) for $L$ implies (M2) (resp. (M4)) for $L$. If #L=2\#L=2, then
$U(\sigma)=\#\{\tau\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}$ $=\#\{\tau\notin L\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}+\#\{\tau\in L\mid\sigma\overset{\mathrm{nc}}{\prec}\tau\}$
@@
$D(\sigma)=\#\{\nu\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}$ $=\#\{\nu\notin L\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}+\#\{\nu\in L\mid\nu\overset{\mathrm{nc}}{\prec}\sigma\}$
@@
and so (M2) and (M4) hold, because $\sigma$ is weakly critical. This means that $f$ is discrete Morse. ∎
Corollary 3.3.
Every Morse function $f$ is a discrete Morse-Bott function such that, for any reduced collection $C\in\mathcal{C}_{f}$, $C\leq 2$ and $C=L(C)$.
Proof.
By Theorem 3.2, we have that $C\leq\#L(C)\leq 2$ for any reduced collection $C\in\mathcal{C}_{f}$. If #C=2\#C=2, then #L(C)=2\#L(C)=2 and so $C=L(C)$. Thus, we may assume that #C=1\#C=1. If #L(C)>1\#L(C)>1, then Theorem 3.2 implies that #L(C)=2\#L(C)=2 and $C=L(C)$, which contradicts the assumption #C=1\#C=1. Therefore, #L(C)=1\#L(C)=1 and so $C=L(C)$. ∎